# Communication-Efficient and Privacy-Adaptable Mechanism for Federated Learning

### v1

@@ Abstract
Training machine learning models on decentralized private data via federated learning (FL) poses two key challenges: communication efficiency and privacy protection. In this work, we address these challenges within the trusted aggregator model by introducing a novel approach called the Communication-Efficient and Privacy-Adaptable Mechanism (CEPAM), achieving both objectives simultaneously. In particular, CEPAM leverages the rejection-sampled universal quantizer (RSUQ), a construction of randomized vector quantizer whose resulting distortion is equivalent to a prescribed noise, such as Gaussian or Laplace noise, enabling joint differential privacy and compression. Moreover, we analyze the trade-offs among user privacy, global utility, and transmission rate of CEPAM by defining appropriate metrics for FL with differential privacy and compression. Our CEPAM provides the additional benefit of privacy adaptability, allowing clients and the server to customize privacy protection based on required accuracy and protection. We assess CEPAM’s utility performance using MNIST dataset, demonstrating that CEPAM surpasses baseline models in terms of learning accuracy.
I Introduction
@@
An alternative approach relies on randomized quantization, such as subtractive dithering [14, 15, 16, 17, 18]. Previous works [19, 20, 21, 22, 23] have employed this method to develop a joint privacy and quantization mechanism for FL. However, these approaches are limited to scalar quantization, which is less effective than vector quantization in terms of compression ratio. Moreover, Lang et al. [24] investigated the use of subtractive dithering to generate privacy-preserving noises such as multivariate $t$-distributed and Laplace noises. However, their proposed mechanism did not encompass the Gaussian mechanism.
@@
Channel simulation [25, 26, 27, 28, 29, 30] is technique that can achieve the same goal. Recognizing that the Gaussian distribution can be represented as a mixture of uniform distributions [31, 32], Agustsson et al. [33] combined subtractive dithering with layered construction [34] to simulate 1D Gaussian channel. This approach further developed in [35] to simulate any 1D unimodal additive noise distribution using randomized scaling and offsets. Recently, a new construction in [36] named RSUQ, which is a randomized vector quantization, was introduced to exactly simulate any multivariate nonuniform continuous additive noise distribution with finite communication by extending 1D methods to multivariate distributions. However, [36] did not consider FL setup. A closely related DP mechanism compressor, based on Poisson functional representation [27], was observed in [37].
@@
In this work, we propose a joint mechanism, CEPAM, aimed at enhancing privacy while compressing model updates in FL. Our proposed CEPAM leverages RSUQ [36], building upon its capability to convert quantization distortion into an additive noise term with adjustable variance, independent of the quantized data. Within CEPAM, both clients and the server can tailor the privacy protection mechanism, offering the choice between the exact Gaussian mechanism (CEPAM-Gaussian) or the exact Laplace mechanism (CEPAM-Laplace), based on the required accuracy level and privacy protection. Additionally, the inherent nature of RSUQ as a vector quantizer gives CEPAM the advantage in reducing the compression ratio compared to its scalar counterparts [22, 23, 33, 34, 35].
@@
We theoretically analyze the performance of CEPAM by studying trade-offs between privacy, rate and accuracy. Specifically, we examine the FL convergence of stochastic gradient descent (SGD)-based federated averaging (FedAvg) with CEPAM in a statistically heterogeneous setup [1, 4], where the dataset available at each user follows different distributions. Additionally, we analyze the privacy of CEPAM-Gaussian and CEPAM-Laplace using techniques from privacy amplification [38, 39]. We validate our theoretical results by comparing CEPAM against other baselines using the MNIST dataset with MLP and CNN architectures. Experimental results demonstrate that CEPAM-Gaussian achieves a 0.5-2.0% improvement in accuracy compared to other baselines.
I-A Related Works
While the original FedAvg [1] works even when the data are non-iid, Stich [40] proposed Local SGD by assuming that the data is iid and provided a theoretical analysis. Later, Li et al. [41] extended Stich’s Local SGD to the heterogeneous data setup and provided a theoretical analysis. Moreover, federated learning with quantization was considered in [8, 9].
@@ I-A Related Works
Write $H(X)$ for the entropy and in bits. Logarithms are to the base $2$. For $\mathcal{A},\mathcal{B}\subseteq\mathbb{R}^{n}$, $\beta\in\mathbb{R}$, $\mathbf{G}\in\mathbb{R}^{n\times n}$, $\mathbf{x}\in\mathbb{R}^{n}$ write $\beta\mathcal{A}:=\{\beta\mathbf{z}:\,\mathbf{z}\in\mathcal{A}\}$, $\mathbf{G}\mathcal{A}:=\{\mathbf{G}\mathbf{z}:\,\mathbf{z}\in\mathcal{A}\}$, $\mathcal{A}+\mathbf{x}:=\{\mathbf{z}+\mathbf{x}:\,\mathbf{z}\in\mathcal{A}\}$, $\mathcal{A}+\mathcal{B}=\{\mathbf{y}+\mathbf{z}:\,\mathbf{y}\in\mathcal{A},\,\mathbf{z}\in\mathcal{B}\}$ for the Minkowski sum, $\mathcal{A}-\mathcal{B}=\{\mathbf{y}-\mathbf{z}:\,\mathbf{y}\in\mathcal{A},\,\mathbf{z}\in\mathcal{B}\}$, and $\mu(\mathcal{A})$ for the Lebesgue measure of $\mathcal{A}$. Let $B^{n}:=\{\mathbf{x}\in\mathbb{R}^{n}:\,\|\mathbf{x}\|\leq 1\}$ be the unit $n$-ball.
II System Model and Preliminaries
In this section, we describe the setup of FL framework. We begin by reviewing the conventional FL framework with bit-constrained model updates in Section II-A, followed by an exploration of LRSUQ in Section II-B, and a discussion of differential privacy and privacy amplification in Section II-C. Next, we formulate the problem and identify the desired properties used in the FL framework in Section II-D.
II-A Federated Learning (FL)
@@
$F_{k}(\mathbf{W}):=\frac{1}{n_{k}}\sum_{j=1}^{n_{k}}\ell(\mathbf{W};\xi_{k,j}),$ (2)
@@
$\mathbf{W}_{t+t^{\prime}}^{k}\leftarrow\mathbf{W}_{t}^{k}-\eta_{t+t^{\prime}}\nabla F_{k}^{j_{t}^{k}}(\mathbf{W}_{t}^{k}),\;\;t^{\prime}=1,\ldots,\tau,$ (3)
where $\eta_{t+t^{\prime}}$ is the learning rate, $\nabla F_{k}^{j}(\mathbf{W}):=\nabla F_{k}(\mathbf{W};\xi_{k,j})$ is the gradient computed at a single sample of index $j$, and $j_{t}^{k}$ is the sample index chosen uniformly from the local data $\mathcal{D}^{(k)}$ of client $k$ at time $t$222 In this work, our focus is on analyzing the computation of a single stochastic gradient at each client during every time instance. The FL convergence rates can potentially be enhanced by incorporating mini-batching techniques [40], leaving the detailed analysis for future work. Finally, by assuming that all clients participate in each FL round for simplicity, the server aggregates the $K$ local model updates $\{\mathbf{X}_{t+\tau}^{k}=\mathbf{W}_{t+\tau}^{k}-\mathbf{W}_{t}\}_{k\in\mathcal{K}}$ and computes the new global parameter vector:
@@
By an abuse of notation, suppose the data in client $k$ is iid sampled from the distribution $\mathcal{D}^{(k)}$. Consequently, the overall distribution becomes a mixture of all local distributions: $\mathcal{D}=\sum_{k\in\mathcal{K}}p_{k}\mathcal{D}^{(k)}$. Previous works typically assume that the data is iid generated by or partitioned among the $K$ clients, i.e., for all $k\in\mathcal{K}$, $\mathcal{D}^{(k)}=\mathcal{D}$. However, we consider a scenario where the data is non-iid (or heterogeneous), implying that $F_{k}$ could potentially be an arbitrarily poor approximation to $F$.
@@
II-B Rejection-Sampled Universal Quantizer (RSUQ)
@@
The quantization error incurred from the aforementioned quantization schemes approximately follows a uniform distribution over a basic cell of a lattice. It may be desirable to have the quantization error follow a uniform distribution over an arbitrary set, rather than being distributed uniformly over a basic cell. RSUQ is a randomized quantizer where the quantization error is uniformly distributed over a set $\mathcal{A}$, a subset of a basic cell. This quantization scheme is based on applying rejection sampling on top of SDQ. Intuitively, we keep generating new dither signals until the quantization error falls in $\mathcal{A}$.
Definition 2
[36, Definition 4] Given a basic cell $\mathcal{P}$ of the lattice $\mathbf{G}\mathbb{Z}^{n}$, a subset $\mathcal{A}\subseteq\mathcal{P}$, and a sequence $S=(\mathbf{V}_{i})_{i\in\mathbb{N}^{+}}$, $\mathbf{V}_{1},\mathbf{V}_{2},\ldots\stackrel{{\scriptstyle iid}}{{\sim}}\mathrm{Unif}(\mathcal{P})$ are i.i.d. dithers, the rejection-sampled universal quantizer (RSUQ) $Q_{\mathcal{A},\mathcal{P}}:\mathbb{R}^{n}\times\prod_{i\in\mathbb{N}^{+}}\mathcal{P}_{i}\to\mathbb{R}^{n}$ for $\mathcal{A}$ against $\mathcal{P}$ is given by
@@
Note that SDQ is a special case of RSUQ where $\mathcal{A}=\mathcal{P}$. It is easy to check that the quantization error is uniform over $\mathcal{A}$ by using the standard rejection sampling argument and the “crypto” lemma [49, Lemma 4.1.1 and Theorem 4.1.1].
Proposition 3
[36, Proposition 5] For any $\mathbf{x}\in\mathbb{R}^{n}$, the quantization error $\mathbf{Z}:=Q_{\mathcal{A},\mathcal{P}}(\mathbf{x},S)-\mathbf{x}$ of RSUQ $Q_{\mathcal{A},\mathcal{P}}$, where $S\sim P_{S}$, follows the uniform distribution over the set $\mathcal{A}$, i.e., $\mathbf{Z}\sim\mathrm{Unif}(\mathcal{A})$.
By using layered construction as in [35, 34], RSUQ can be generalized to simulate an additive noise channel with noise following a continuous distribution.
Definition 4
[36, Definition 12] Given a basic cell $\mathcal{P}$ of the lattice $\mathbf{G}\mathbb{Z}^{n}$, a probability density function $f:\mathbb{R}^{n}\to[0,\infty)$ where $L_{u}^{+}(f)$ is always bounded for $u>0$, and $\beta:(0,\infty)\to[0,\infty)$ satisfying $L_{u}^{+}(f)\subseteq\beta(u)\mathcal{P}$ for $u>0$, and a random pair $S=(U,(\mathbf{V}_{i})_{i\in\mathbb{N}^{+}})$ where the latent variable $U\sim f_{U}$ with $f_{U}(u):=\mu(L_{u}^{+}(f))$, and $\mathbf{V}_{1},\mathbf{V}_{2},\ldots\stackrel{{\scriptstyle iid}}{{\sim}}\mathrm{Unif}(\mathcal{P})$ is a sequence of i.i.d. dither signals, the layered rejection-sampled universal quantizer (LRSUQ) $Q_{f,\mathcal{P}}:\mathbb{R}^{n}\times\mathbb{R}\times\prod_{i\in\mathbb{N}^{+}}\mathcal{P}_{i}\to\mathbb{R}^{n}$ for $f$ against $\mathcal{P}$ is given by
@@
Proposition 5
[36, Proposition 13] Consider LRSUQ $Q_{f,\mathcal{P}}$. For any random input $\mathbf{X}$, the quantization error defined by $\mathbf{Z}:=Q_{f,\mathcal{P}}(\mathbf{X},U,(\mathbf{V}_{i})_{i})-\mathbf{X}$, follows the pdf $f$, independent of $\mathbf{X}$.
II-C Differential Privacy (DP)
Since clients do not directly transmit their data to the server, the original FL framework [1] provides a certain level of privacy. Nonetheless, a significant amount of information can still be inferred from the shared data, e.g., model parameters induced by gradient descent, by potential eavesdroppers within the FL network. Consequently, a privacy mechanism, such as DP, is essential to protect the shared information.
In this section, we delve into a notion of DP [10, 11]. In DP, clients place trust in the server (or data curator) responsible for collecting and holding their individual data in a database $X\in\mathcal{X}$, where $\mathcal{X}$ denotes the collection of databases. The server then introduces privacy-preserving noise to the original datasets or query results through a randomized mechanism $\mathcal{A}$, producing an output $Y=\mathcal{A}(X)\in\mathcal{Y}$, where $\mathcal{Y}$ denotes the set of possible outputs, before sharing them with untrusted data analysts. While this model requires a higher level of trust compared to the local model, it enables the design of significantly more accurate algorithms. Two databases $X$ and $X^{\prime}$ are considered adjacent if they differ in only one entry. More generally, we can define a symmetric adjacent relation $\mathcal{R}\subseteq\mathcal{X}^{2}$ and say that $X$ and $X^{\prime}$ are adjacent databases if $(X,X^{\prime})\in\mathcal{R}$. Here, we review the definition of $(\epsilon,\delta)$-differentially private, initially introduced by Dwork et al. [10].
Definition 6 ($(\epsilon,\delta)$-differential privacy [10])

### v2

@@ Abstract
Training machine learning models on decentralized private data via federated learning (FL) poses two key challenges: communication efficiency and privacy protection. In this work, we address these challenges within the trusted aggregator model by introducing a novel approach called the Communication-Efficient and Privacy-Adaptable Mechanism (CEPAM), achieving both objectives simultaneously. In particular, CEPAM leverages the rejection-sampled universal quantizer (RSUQ), a construction of randomized vector quantizer whose resulting distortion is equivalent to a prescribed noise, such as Gaussian or Laplace noise, enabling joint differential privacy and compression. Our CEPAM provides the additional benefit of privacy adaptability, allowing clients and the server to customize privacy protection based on required accuracy and protection. We theoretically analyze the privacy guarantee of CEPAM and investigate the trade-offs among user privacy and accuracy of CEPAM through experimental evaluations. Moreover, we assess CEPAM’s utility performance using MNIST dataset, demonstrating that CEPAM surpasses baseline models in terms of learning accuracy.
I Introduction and Motivation
@@
An alternative approach relies on randomized quantization, such as subtractive dithering [14, 15, 16, 17, 18]. Previous works [19, 20, 21, 22, 23] have employed this method to develop a joint privacy and quantization mechanism for FL. However, these approaches are limited to scalar quantization, which is less effective than vector quantization in terms of compression ratio. Moreover, Lang et al. [24] investigated the use of subtractive dithering to generate privacy-preserving noises such as multivariate $t$-distributed and Laplace noises for local DP. However, their proposed mechanism did not encompass the Gaussian mechanism. [^1] The mechanism proposed in [24] adds a carefully designed, low-power privacy-preserving noise to the model update before the dithered quantization step. However, the authors do not specify how to design the requisite infinitesimal noise to generate the Gaussian noise for local DP, a task that is highly non-trivial. In contrast, our proposed CEPAM specifies how to generate Gaussian noise for central DP.
@@
Channel simulation [25, 26, 27, 28, 29, 30, 31, 32] is technique that can achieve the same goal. Recognizing that the Gaussian distribution can be represented as a mixture of uniform distributions [33, 34], Agustsson et al. [35] combined subtractive dithering with layered construction [36] to simulate 1D Gaussian channel. This approach was further developed in [37] to simulate any 1D unimodal additive noise distribution using randomized scaling and offsets. Recently, a new construction in [38] named RSUQ, which is a randomized vector quantization, was introduced to exactly simulate any multivariate nonuniform continuous additive noise distribution with finite communication by extending 1D methods to multivariate distributions. However, [38] did not consider FL setup.
@@
In this work, we propose a joint mechanism, CEPAM, aimed at enhancing privacy while compressing model updates in FL. Our proposed CEPAM leverages RSUQ [38], building upon its capability to convert quantization distortion into an additive noise term with adjustable variance, independent of the quantized data. Within CEPAM, both clients and the server can tailor the privacy protection mechanism, offering the choice between the exact Gaussian mechanism (CEPAM-Gaussian) or Laplace mechanism (CEPAM-Laplace), based on the required accuracy level and privacy protection. Additionally, the inherent nature of RSUQ as a vector quantizer gives CEPAM the advantage in reducing the compression ratio compared to its scalar counterparts [21, 22, 35, 36, 37].
@@
We theoretically analyze the privacy guarantee of CEPAM and investigate the trade-offs among user privacy and accuracy of CEPAM through experimental evaluations. Additionally, we analyze the privacy of CEPAM-Gaussian and CEPAM-Laplace using techniques from privacy amplification [39, 40]. We validate our theoretical results by comparing CEPAM-Gaussian against other baselines using the MNIST dataset with MLP and CNN architectures. The experimental results demonstrate that for MLP and CNN models, CEPAM-Gaussian achieves improvements of 0.6-2.0% and 0.4-1.0% in accuracy, respectively, compared to other baselines.
The rest of this paper is organized as follows. Section II summarizes recent related works and compares them to our approach. Section III reviews necessary preliminaries, presents the system model, and identifies the requirements for joint quantization and privacy in FL. Section IV details the proposed joint quantization and privacy scheme, and Section V provides its theoretical performance analysis. Finally, experimental results are presented in Section VI.
II Related Works
The channel simulation problem [25, 41, 42] is closely related to Shannon’s lossy compression problem. In this setup [25], an encoder compresses an input source $X$ into a message $M$, which a decoder then uses to generate an output $M$. The encoder and decoder may share common randomness. The objective is to ensure that $Y$ follows a prescribed conditional distribution given $X$. Channel simulation was initially studied in the asymptotic setting [25, 41, 42], where $X$ is a sequence, and has also been investigated in the one-shot setting [26, 43, 27, 44], where $X$ is a scalar. RSUQ [38] can be regarded as a one-shot channel simulation scheme for simulating an additive noise channel with an arbitrary continuous noise distribution. For a comprehensive treatment of the subject, we direct the interested reader to the monograph in [32].
Federated learning (FL) was initially introduced by McMahan et al. [1] as a collaborative learning approach that allows model training without the need to gather user-specific data. Research in FL focuses mainly on improving communication efficiency [1, 5, 8, 45, 46, 47, 9, 48, 49], ensuring data privacy [50, 51], or an approach that addresses both aspects [21, 22, 31]. This work is focused on an approach that addresses the issues of communication efficiency and data privacy simultaneously.
The communication cost in federated learning can be mitigated through two primary methods: reducing the number of communication rounds or compressing model updates (e.g., parameter vectors or stochastic gradients) before transmission. While the original FedAvg algorithm [1] operates heuristically under heterogeneous data (i.e., non-iid data) and device inactivity (i.e., stragglers), Local SGD explicitly reduces communication frequency by performing multiple local update steps before synchronization. However, the convergence guarantees for Local SGD, which were first established by Stich [47] for strongly-convex and smooth objectives, relied on the idealized assumptions of iid data and no stragglers. Li et al. [49] later extended this analysis by providing theoretical guarantees for Local SGD in settings with heterogeneous data and inactive devices. Subsequently, Shlezinger et al. [9] integrated compression into this framework with UVeQFed, which incorporates quantization for heterogeneous settings.
Another major method to reducing communication costs is the compression of stochastic gradients prior to transmission [8, 46, 9]. This includes methods like quantization, which constrains the number of bits used per value. A key technique is unbiased quantization [8], which ensures the compressed gradients remain unbiased estimators of their true values. In contrast, sparsification techniques [45, 46] improve communication efficiency by reducing the number of non-zero entries in the stochastic gradients. Building on these advances, we propose a novel Local SGD variant that unifies quantization with a privacy-preserving mechanism based on RSUQ, simultaneously addressing communication efficiency and data confidentiality within a single framework.
A recent trend involves developing approaches that jointly address communication efficiency and privacy [19, 29, 21, 22, 24, 23, 31, 52]. For central DP, [21, 22] used subtractive dithering to achieve this for scalar entries. CEPAM generalizes the scalar counterparts in [21, 22], as it uses RSUQ, which is a vector quantizer. For local DP, [24] proposed injecting low-power noise before dithered quantization step, proving guarantees for Laplace distribution and multidimensional $t$-distribution. However, the authors do not specify how to generate the requisite infinitesimal noise to produce Gaussian noise. In contrast, our CEPAM directly specifies the generation of this Gaussian noise for central DP. In parallel, techniques from channel simulation have inspired other DP compressors [29, 31].
III System Model and Preliminaries
In this section, we describe the setup of FL framework. We begin by reviewing the conventional FL framework with bit-constrained model updates in Section III-A, followed by an exploration of LRSUQ in Section III-B, and a discussion of differential privacy and privacy amplification in Section III-C. Next, we formulate the problem and identify the desired properties used in the FL framework in Section III-D.
@@ III System Model and Preliminaries
Write $H(X)$ for the entropy and in bits. Logarithms are to the base $2$. For $\mathcal{A},\mathcal{B}\subseteq\mathbb{R}^{n}$, $\beta\in\mathbb{R}$, $\mathbf{G}\in\mathbb{R}^{n\times n}$, $\mathbf{x}\in\mathbb{R}^{n}$ write $\beta\mathcal{A}:=\{\beta\mathbf{z}:\,\mathbf{z}\in\mathcal{A}\}$, $\mathbf{G}\mathcal{A}:=\{\mathbf{G}\mathbf{z}:\,\mathbf{z}\in\mathcal{A}\}$, $\mathcal{A}+\mathbf{x}:=\{\mathbf{z}+\mathbf{x}:\,\mathbf{z}\in\mathcal{A}\}$, $\mathcal{A}+\mathcal{B}=\{\mathbf{y}+\mathbf{z}:\,\mathbf{y}\in\mathcal{A},\,\mathbf{z}\in\mathcal{B}\}$ for the Minkowski sum, $\mathcal{A}-\mathcal{B}=\{\mathbf{y}-\mathbf{z}:\,\mathbf{y}\in\mathcal{A},\,\mathbf{z}\in\mathcal{B}\}$, and $\mu(\mathcal{A})$ for the Lebesgue measure of $\mathcal{A}$. Let $B^{n}:=\{\mathbf{x}\in\mathbb{R}^{n}:\,\|\mathbf{x}\|\leq 1\}$ be the unit $n$-ball. For a function $f:\mathbb{R}^{n}\rightarrow\mathbb{R}$, its superlevel set is defined as $L_{u}^{+}(f):=\{\mathbf{x}\in\mathbf{R}^{n}:f(\mathbf{x})\geq u\}$.
III-A Federated Learning (FL)
@@
$F_{k}(\mathbf{W}){\color[rgb]{0,0,0}\equiv F_{k}(\mathbf{W},\mathcal{D}^{(k)})}:=\frac{1}{n_{k}}\sum_{j=1}^{n_{k}}\ell(\mathbf{W};\xi_{k,j}),$ (2)
@@
$\mathbf{W}_{t+t^{\prime}}^{k}\leftarrow\mathbf{W}_{t{\color[rgb]{0,0,0}+t^{\prime}-1}}^{k}-\eta_{t+t^{\prime}}\nabla F_{k}^{j_{t{\color[rgb]{0,0,0}+t^{\prime}-1}}^{k}}(\mathbf{W}_{t{\color[rgb]{0,0,0}+t^{\prime}-1}}^{k}),\;\;t^{\prime}=1,\ldots,\tau,$ (3)
where $\eta_{t+t^{\prime}}$ is the learning rate, $\nabla F_{k}^{j}(\mathbf{W}):=\nabla F_{k}(\mathbf{W};\xi_{k,j})$ is the gradient computed at a single sample of index $j$, and $j_{t}^{k}$ is the sample index chosen uniformly from the local data $\mathcal{D}^{(k)}$ of client $k$ at time tt. [^3] In this work, our focus is on analyzing the computation of a single stochastic gradient at each client during every time instance. The FL convergence rates can potentially be enhanced by incorporating mini-batching technique [47], leaving the detailed analysis for future work. Finally, by assuming that all clients participate in each FL round for simplicity, the server aggregates the $K$ local model updates $\{\mathbf{X}_{t+\tau}^{k}=\mathbf{W}_{t+\tau}^{k}-\mathbf{W}_{t}\}_{k\in\mathcal{K}}$ and computes the new global parameter vector:
@@
The dataset $\mathcal{D}^{(k)}$ inherently induces a distribution. By an abuse of notation, we also denote this induced distribution as $\mathcal{D}^{(k)}$. Suppose the data in client $k$ is iid sampled from the induced distribution $\mathcal{D}^{(k)}$. Consequently, the overall distribution becomes a mixture of all local distributions: $\mathcal{D}=\sum_{k\in\mathcal{K}}p_{k}\mathcal{D}^{(k)}$. Previous works typically assume that the data is iid generated by or partitioned among the $K$ clients, i.e., for all $k\in\mathcal{K}$, $\mathcal{D}^{(k)}=\mathcal{D}$. However, we consider a scenario where the data is non-iid (or heterogeneous), implying that $F_{k}$ could potentially be an arbitrarily poor approximation to $F$.
@@
III-B Rejection-Sampled Universal Quantizer (RSUQ)
@@
The quantization error incurred from the aforementioned quantization schemes approximately follows a uniform distribution over a basic cell of a lattice. It may be desirable to have the quantization error follow a uniform distribution over an arbitrary set, rather than being distributed uniformly over a basic cell. RSUQ is a randomized quantizer where the quantization error is uniformly distributed over a set $\mathcal{A}$, a subset of a basic cell. This quantization scheme is based on applying rejection sampling on top of SDQ. Intuitively, we keep generating new dither signals until the quantization error falls in $\mathcal{A}$. [^4] It is easy to see that the acceptance probability is $\mu(\mathcal{A})/\mu(\mathcal{P})$.
Definition 2.
[38, Definition 4] Given a basic cell $\mathcal{P}$ of the lattice $\mathbf{G}\mathbb{Z}^{n}$, a subset $\mathcal{A}\subseteq\mathcal{P}$, and a sequence $S=(\mathbf{V}_{i})_{i\in\mathbb{N}^{+}}$, $\mathbf{V}_{1},\mathbf{V}_{2},\ldots\stackrel{{\scriptstyle iid}}{{\sim}}\mathrm{Unif}(\mathcal{P})$ are i.i.d. dithers, the rejection-sampled universal quantizer (RSUQ) $Q_{\mathcal{A},\mathcal{P}}:\mathbb{R}^{n}\times\prod_{i\in\mathbb{N}^{+}}\mathcal{P}_{i}\to\mathbb{R}^{n}$ for $\mathcal{A}$ against $\mathcal{P}$ is given by
@@
Note that SDQ is a special case of RSUQ where $\mathcal{A}=\mathcal{P}$. It is easy to check that the quantization error is uniform over $\mathcal{A}$ by using the standard rejection sampling argument and the “crypto” lemma [60, Lemma 4.1.1 and Theorem 4.1.1].
Proposition 3.
[38, Proposition 5] For any $\mathbf{x}\in\mathbb{R}^{n}$, the quantization error $\mathbf{Z}:=Q_{\mathcal{A},\mathcal{P}}(\mathbf{x},S)-\mathbf{x}$ of RSUQ $Q_{\mathcal{A},\mathcal{P}}$, where $S\sim P_{S}$, follows the uniform distribution over the set $\mathcal{A}$, i.e., $\mathbf{Z}\sim\mathrm{Unif}(\mathcal{A})$.
By using layered construction as in [37, 36], RSUQ can be generalized to simulate an additive noise channel with noise following a continuous distribution. Consider a pdf $f:\mathbb{R}^{n}\to[0,\infty)$ and write its superlevel set as
$L_{u}^{+}(f)=\{\mathbf{z}\in\mathbb{R}^{n}:\,f(\mathbf{z})\geq u\}.$
Let $f_{U}(u):=\mu(L_{u}^{+}(f))$ for $u>0$, which is also a pdf. If we generate $U\sim f_{U}$, and then $\mathbf{Z}|\{U=u\}\sim\mathrm{Unif}(L_{u}^{+}(f))$, then we have $\mathbf{Z}\sim f$ by the fundamental theorem of simulation [64].
Definition 4.
[38, Definition 12] Given a basic cell $\mathcal{P}$ of the lattice $\mathbf{G}\mathbb{Z}^{n}$, a probability density function $f:\mathbb{R}^{n}\to[0,\infty)$ where $L_{u}^{+}(f)$ is always bounded for $u>0$, and $\beta:(0,\infty)\to[0,\infty)$ satisfying $L_{u}^{+}(f)\subseteq\beta(u)\mathcal{P}$ for $u>0$, and a random pair $S=(U,(\mathbf{V}_{i})_{i\in\mathbb{N}^{+}})$ where the latent variable $U\sim f_{U}$ with $f_{U}(u):=\mu(L_{u}^{+}(f))$, and $\mathbf{V}_{1},\mathbf{V}_{2},\ldots\stackrel{{\scriptstyle iid}}{{\sim}}\mathrm{Unif}(\mathcal{P})$ is a sequence of i.i.d. dither signals, the layered rejection-sampled universal quantizer (LRSUQ) $Q_{f,\mathcal{P}}:\mathbb{R}^{n}\times\mathbb{R}\times\prod_{i\in\mathbb{N}^{+}}\mathcal{P}_{i}\to\mathbb{R}^{n}$ for $f$ against $\mathcal{P}$ is given by
@@
Proposition 5.
[38, Proposition 13] Consider LRSUQ $Q_{f,\mathcal{P}}$. For any random input $\mathbf{X}$, the quantization error defined by $\mathbf{Z}:=Q_{f,\mathcal{P}}(\mathbf{X},U,(\mathbf{V}_{i})_{i})-\mathbf{X}$, follows the pdf $f$, independent of $\mathbf{X}$.
III-C Differential Privacy (DP)
Since clients do not directly transmit their data to the server, the original FL framework [1] provides a certain level of privacy. Nonetheless, a significant amount of information can still be inferred from the shared data, e.g., model parameters induced by gradient descent, by potential eavesdroppers within the FL network [65, 66, 67]. Consequently, a privacy mechanism, such as DP, is essential to protect the shared information.
In this section, we delve into a notion of (central) DP [10, 11]. [^5] Herein, DP refers to central DP. In DP, clients place trust in the server (or data curator) responsible for collecting and holding their individual data in a database $X\in\mathcal{X}$, where $\mathcal{X}$ denotes the collection of databases. The server then introduces privacy-preserving noise to the original datasets or query results through a randomized mechanism $\mathcal{F}$, producing an output $Y=\mathcal{F}(X)\in\mathcal{Y}$, where $\mathcal{Y}$ denotes the set of possible outputs, before sharing them with untrusted data analysts. While this model requires a higher level of trust compared to the local model, it enables the design of significantly more accurate algorithms. Two databases $X$ and $X^{\prime}$ are considered adjacent if they differ in only one entry. More generally, we can define a symmetric adjacent relation $\mathcal{R}\subseteq\mathcal{X}^{2}$ and say that $X$ and $X^{\prime}$ are adjacent databases if $(X,X^{\prime})\in\mathcal{R}$. Here, we review the definition of $(\epsilon,\delta)$-differentially private, initially introduced by Dwork et al. [10].
Definition 6 ($(\epsilon,\delta)$-differential privacy [10]).